# Polylogarithmic motivic Chabauty–Kim for $\mathbb{P}^{1}\setminus\{0,1,\infty\}$: the geometric step via resultants

### v2

@@ 4.3. The geometric step in coordinates
defined in Definition 4.1, a problem in linear algebra.
@@ 5. An upper bound algorithm
defined in Definition 4.1, where
@@
  • each $\Phi^{\sigma_{2n+1}}_{e_{1}e_{0}^{2n}}$ is assigned degree $12n+$ 1.
@@
We call an element of $\ker(\theta^{\#}_{d,v})$ a polylogarithmic motivic Chabauty–Kim function of depth $d$ and degree $v$.
@@ 5.1. Existence of Chabauty–Kim functions in high depth
Fix a depth $d$ and a degree $v$ in the sense of Definition 5.1. Both $(\mathcal{O}(U_{S})_{\leq d}[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}])_{v}$ and $(R_{\Phi,d})_{v}$ have natural monomial bases, so we can view $\theta_{d,v}^{\#}$ as a matrix $M(\theta_{d,v}^{\#})$ with coefficients in $\mathcal{O}(U_{S})_{\leq d}$ with respect to these bases. Define
@@
Then, if $p(n;T)$ denotes the number of partitions of an integer $n\in\mathbb{N}$ with respect to a tuple of integers $T$, we have
@@
The numerology stems from the fact that $\Phi_{e_{0}}^{\tau_{\ell}}$ and $\Phi_{e_{1}}^{\tau_{\ell}}$ all have degree $1$ for $\ell\in S$, while $\Phi^{\sigma_{2k+1}}_{e_{1}e_{0}^{2k}}$ has degree $12k+$ 1.
@@
If we fix $d$ to be a large enough constant so that $d>2|S|+(d-1)/2$, then for $v$ large enough we have $\dim_{\operatorname{PL}}(d,v)>\dim_{\Phi}(d,v)$. ∎
@@
On the one hand, $\dim_{\operatorname{PL}}(4,v)$ counts the number of ways to express $v$ as a sum of distinct weights of the form $(1,1,2,3,4)$, with repetitions and without order, while on the other hand $\dim_{\Phi}(3,v)$ counts the number of ways to express $v$ as a sum of distinct weights of the form $(1,1,1,1,3)$, under similar constraints. A $\operatorname{PL}$-partition of $v$ is a vector $(a,b,c,d,e)$ of non-negative integers, having the property that $a+b+2c+3d+4e=v$. One way to prove the inequality $\dim_{\operatorname{PL}}(4,v)\leq D_{\Phi}(3,v)$ is to produce an injection $(a,b,c,d,e)\mapsto(a^{\prime},b^{\prime},c^{\prime},d^{\prime},e^{\prime})$ from the set of $\operatorname{PL}$-partitions of an arbitrary fixed $v$ in depth $4$ to the set of $\Phi$-partitions of the same $v$ in depth $3$. An example of such an injection is given by mapping a vector $(a,b,c,d,e)$ to $(a+\Delta(a,b,c,d,e),b,c,d,e)$, where
@@
when $d=2$ and $v=2$. This phenonomenon explains the relative ease of solving the geometric step when $|S|=1$ (see Section 5.2.1). When $|S|=2$, the following table exhibits the first degree $v$ at which the inequality (5.2) becomes true in depth $d$ for $1\leq d\leq 30$:
@@
The blank rows in depths $d=1,2,3,4,5$ for $|S|=2$ indicate that the inequality $\dim_{\operatorname{PL}}(d,v)>\dim_{\Phi}(d,v)$ is never satisfied, in agreement with Proposition 5.3.
@@ 5.2. A naive approach: computing the kernel of $\theta^{\#}_{d,v}$
The matrix $M(\theta^{\#}_{d,v})$ has size $\dim_{\Phi}(d,v)\times\dim_{\operatorname{PL}}(d,v)$ with coefficients in $\mathcal{O}(U_{S})_{\leq d}$ and by Proposition 5.2, once $d$ and $v$ are large enough, it is guaranteed to have a non-trivial kernel. Hence, one could attempt to find Chabauty–Kim functions by explicitly computing these matrices.
@@
One could attempt a similar strategy in the general case, however, it quickly proves computationally infeasible. The matrix $M(\theta^{\#}_{d,v})$ has coefficients in the shuffle algebra $\mathcal{O}(U_{S})_{\leq d}$. As an abstract algebra, the shuffle algebra $\mathcal{O}(U_{S})$ is isomorphic to the free commutative algebra $\mathbb{Q}[\mathcal{L}]$ generated by the set $\mathcal{L}$ of Lyndon words of $\Sigma$ over $\mathbb{Q}$ [LYNDON, §4 Thm. (i)]. We may thus view $M(\theta^{\#}_{d,v})$ as a matrix with coefficients in a polynomial ring over $\mathbb{Q}$.
When $|S|=2$, by Table 1, the minimal matrix size at which (5.2) holds is for $(d,v)=(14,65)$. When $d=14$, $296$ distinct Lyndon words show up in the image of $\theta^{\#}_{d}$. Thus, this brute force approach requires computing the kernel of a matrix of size $7354311\times 7370562$ over the ring $\mathbb{Q}[X_{1},\ldots,X_{296}]$, which is infeasible.
A key contribution of this paper, which we outline in the next section, is a method to produce polylogarithmic motivic Chabauty–Kim functions without computing this kernel using matrix operations. In the remainder of this section, we show how the naive method can nevertheless be salvaged to give an upper bound on the dimension of the kernel of $M(\theta^{\#}_{d,v})$.
@@ 5.3. Non-existence of Chabauty–Kim functions in low degree
Recall that $M(\theta^{\#}_{d,v})$ is a matrix over the shuffle algebra $\mathcal{O}(U_{S})_{\leq d}$, which we can view as a polynomial ring generated by the Lyndon words $\mathcal{L}_{\geq-d}$ of degree $\geq-d$ on the generators $\Sigma$. Write $\mathcal{O}(U_{S})_{\leq d}=\mathbb{Q}[\mathcal{L}_{\geq-d}]=\mathbb{Q}[X_{1},\ldots,X_{n}]$, where $n$ is the number of Lyndon words on the set $\Sigma$ of degree $\geq-d$.
Let $x\in\mathbb{Z}^{n}$. Then, we can evaluate $M(\theta^{\#}_{d,v})$ at $x$ to obtain a matrix $M(\theta^{\#}_{d,v})(x)$ of the same size as $M(\theta^{\#}_{d,v})$, but with coefficients in $\mathbb{Q}$. Observe that if $f\in\ker(\theta^{\#}_{d,v})\subseteq\mathcal{O}(U_{S})_{\leq d}[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}]$, then its evaluation $f(x)\in\mathbb{Q}[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}]$ is in the kernel of $M(\theta^{\#}_{d,v})(x)$.
@@
Indeed, in order for the inequality $\mathrm{rank}_{\mathcal{O}(U_{S})_{\leq d}}\ker M(\theta^{\#}_{d,v})<\dim_{\mathbb{Q}}\ker M(\theta^{\#}_{d,v})(x)$ to hold, the rank $r$ of $M(\theta^{\#}_{d,v})$ must be greater than the rank of $M(\theta^{\#}_{d,v})(x)$. The point $x$ must therefore be a root of all the non-trivial $r\times r$ minors of $M(\theta^{\#}_{d,v})$. Thus, $x$ must lie in the intersection of finitely many hypersurfaces in $\mathbb{A}^{n}$, so in particular, in a Zariski closed subset.
The utility of the observations made in Proposition 5.4 and Remark 5.5 is that, while computing the kernel of $M(\theta^{\#}_{d,v})$ is computationally hard, computing the dimension of the kernel of $M(\theta^{\#}_{d,v})(x)$ for any $x\in\mathbb{Z}^{n}$ is computationally easy.
Motivated by Proposition 5.4, we have implemented the following algorithm in SageMath [sagemath]. It takes as input a tuple of integers $(s,d,v)$, where $s=|S|$, and outputs an integer $r$, which is the dimension of $\ker(M(\theta^{\#}_{d,v})(x))$ for a random $x\in\mathbb{Z}^{n}$. By Proposition 5.4, if the output is 00, then there are no Chabauty–Kim functions in depth $d$ and degree $v$ for $|S|=s$. Moreover, if the output is $r$ for several iterations of the same tuple $(s,d,v)$, then it is likely that the dimension of $M(\theta^{\#}_{d,v})(x)$ is exactly $r$.
@@ 1. $(i)$
    Compute all monomials of degree $v$ in the variables $\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}$. These monomials form a basis of $(\mathcal{O}(U_{S})_{\leq d}[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}])_{v}$ as an $\mathcal{O}(U_{S})_{\leq d}$-module;
@@ 2. $(ii)$
    Compute all monomials of degree $v$ in the variables $\Phi_{\lambda}^{\rho}$ for $\rho\in\{\tau_{\ell},\sigma_{2k+1}\}$ and $\lambda=e_{0},e_{1},e_{1}e_{0}^{2k}$. These monomials form a basis of $(R_{\Phi,d})_{v}$ as an $\mathcal{O}(U_{S})_{\leq d}$-module;
@@ 3. $(iii)$
    Compute the set of Lyndon words $\mathcal{L}_{\geq-d}$ of $\Sigma$ to express $\mathcal{O}(U_{S})_{\leq d}$ as a polynomial algebra $\mathbb{Q}[\mathcal{L}_{\geq-d}]=\mathbb{Q}[X_{1},\ldots,X_{n}]$;
@@ 4. $(iv)$
    Choose a random $n$-tuple $x\in\mathbb{Z}^{n}$;
@@ 5. $(v)$
    Build the matrix $M(\theta_{d,v}^{\#})$ with coefficients in $\mathcal{O}(U_{S})_{\leq d}$;
@@ 6. $(vi)$
    Convert the matrix $M(\theta_{d,v}^{\#})$ into the form given by the isomorphism $\mathcal{O}(U_{S})_{\leq d}\cong\mathbb{Q}[X_{1},\ldots,X_{n}]$, and evaluate the matrix at $x$ to obtain a matrix with coefficients in $\mathbb{Q}$;
@@ 7. $(vii)$
The computation was carried out on The Ohio State University’s high performance cluster, and took around $66$ hours across $96$ cores, using $432$GB of memory. The overwhelming majority of this time was spent converting words in the shuffle algebra $\mathcal{O}(U_{S})$ into the dual PBWL basis (Poincaré–Birkhoff–Witt basis constructed on the Lyndon basis) using an improved version of SageMath’s built-in ShuffleAlgebra.to_dual_pbw_element method, rewritten to enable parallelisation, and to fix inefficiencies.
@@ 6. The geometric step via resultants
Inspired by ideas in [DCJ], we introduce a method using resultants, whose goal is to produce a polylogarithmic motivic Chabauty–Kim function of depth $\leq d$, i.e., a non-zero element in the kernel of the homomorphism (Definition 4.1)
@@
We assume from now on that $|S|=2$ and write $S=\{p,q\}$, so that $\Sigma_{-1}=\{\tau_{p},\tau_{q}\}$. With this assumption, the output of this section is an explicit non-trivial element in $\ker(\theta^{\#}_{6,18})$ (Definition 6.9 and Theorem 6.10).
@@ 6.1. The equations defining the map $\theta^{\#}$
Making Definition 4.1 explicit in the case $|S|=2$ gives the formulas:
@@ 6.2. The elimination process
We now describe the process of “eliminating” variables $\Phi_{\lambda}^{\rho}$, i.e., expressing them in terms of other variables $\Phi_{\lambda^{\prime}}^{\rho^{\prime}}$ and images under $\theta^{\#}$ in order to obtain a Chabauty–Kim function. The system of equations described in Section 6.1 is linear with respect to the variables $\Phi^{\sigma_{2n-1}}_{e_{1}e_{0}^{2n-2}}$ and the variables $\Phi_{e_{1}}^{\tau_{p}}$ and $\Phi_{e_{1}}^{\tau_{q}}$. Consequently, we first seek to eliminate these variables. Eventually, the elimination of the variables $\Phi^{\tau_{p}}_{e_{0}}$ and $\Phi^{\tau_{q}}_{e_{0}}$ requires the use of a resultant. Throughout this section, we will be working with polynomials with coefficients in the fraction field $K=\mathrm{Frac}(\mathcal{O}(U_{S}))$.
@@ 6.2.2. Elimination of $\Phi_{e_{1}}^{\tau_{p}}$ and $\Phi_{e_{1}}^{\tau_{q}}$
$\Delta\Phi_{e_{1}}^{\tau_{p}}=\frac{1}{2}f_{[\tau_{q},\tau_{p}]}\theta^{\#}(\operatorname{Li}_{1}^{\mathfrak{u}})\Phi_{e_{0}}^{\tau_{p}}-f_{\tau_{q}}\theta^{\#}(F^{|1|}_{2,2}),$
@@
Note that $\theta^{\#}(F^{|1|}_{2,2})\neq 0$ even though $F^{|1|}_{2,2}\in\mathcal{I}^{\operatorname{PL}}_{S^{\prime},2}$ when $|S^{\prime}|=1$.

### v3

@@ 4.3. The geometric step in coordinates
defined in Definition˜4.1, a problem in linear algebra.
@@ 5. An upper bound algorithm
defined in Definition˜4.1, where
@@
  • each $\Phi^{\sigma_{2n+1}}_{e_{1}e_{0}^{2n}}$ is assigned degree $2n+1$.
@@
[…]
@@ 5.1. Existence of Chabauty–Kim functions in high depth
Fix a depth $d$ and a degree $v$ in the sense of Definition˜5.1. Both $(\mathcal{O}(U_{S})_{\leq d}[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}])_{v}$ and $(R_{\Phi,d})_{v}$ have natural monomial bases, so we can view $\theta_{d,v}^{\#}$ as a matrix $M(\theta_{d,v}^{\#})$ with coefficients in $\mathcal{O}(U_{S})_{\leq d}$ with respect to these bases. Define
@@
Given a tuple of positive integers $\alpha$ of length $|\alpha|$ and a positive integer $n$, denote the number of partitions of $n$ with respect to $\alpha$ by $p(n;\alpha):=\#\{\beta\in(\mathbb{Z}_{\geq 0})^{|\alpha|}\mid\langle\alpha,\beta\rangle=n\}$. We then have
@@
The numerology stems from the fact that $\Phi_{e_{0}}^{\tau_{\ell}}$ and $\Phi_{e_{1}}^{\tau_{\ell}}$ all have degree $1$ for $\ell\in S$, while $\Phi^{\sigma_{2k+1}}_{e_{1}e_{0}^{2k}}$ has degree $2k+1$.
Remark 5.2.
@@
If we fix $d$ to be a large enough constant so that d>2|S|+(d1)/2d>2|S|+(d-1)/2, then for $v$ large enough we have $\dim_{\operatorname{PL}}(d,v)>\dim_{\Phi}(d,v)$. ∎
@@
On the one hand, $\dim_{\operatorname{PL}}(4,v)$ counts the number of ways to express $v$ as a sum of distinct weights of the form $(1,1,2,3,4)$, with repetitions and without order, while on the other hand $\dim_{\Phi}(3,v)$ counts the number of ways to express $v$ as a sum of distinct weights of the form $(1,1,1,1,3)$, under similar constraints. A $\operatorname{PL}$-partition of $v$ is a vector $(a,b,c,d,e)$ of non-negative integers, having the property that $a+b+2c+3d+4e=v$. One way to prove the inequality $\dim_{\operatorname{PL}}(4,v)\leq\dim_{\Phi}(3,v)$ is to produce an injection $(a,b,c,d,e)\mapsto(a^{\prime},b^{\prime},c^{\prime},d^{\prime},e^{\prime})$ from the set of $\operatorname{PL}$-partitions of an arbitrary fixed $v$ in depth $4$ to the set of $\Phi$-partitions of the same $v$ in depth $3$. An example of such an injection is given by mapping a vector $(a,b,c,d,e)$ to $(a+\Delta(a,b,c,d,e),b,c,d,e)$, where
@@
when $d=2$ and $v=2$. This phenomenon explains the relative ease of solving the geometric step when $|S|=1$ (see Section˜5.2.1). When $|S|=2$, the following table exhibits the first degree $v$ at which the inequality (5.2) becomes true in depth $d$ for $1\leq d\leq 30$:
@@
This entries of this table were calculated using SageMath [sagemath]. The code is available at our Github repository.
The blank rows in depths $d=1,2,3,4,5$ for $|S|=2$ indicate that the inequality $\dim_{\operatorname{PL}}(d,v)>\dim_{\Phi}(d,v)$ is never satisfied, in agreement with Proposition˜5.4. Note that the combination of Remark˜5.2 with the row $d=6$ of Table˜1 implies the existence of a motivic Chabauty–Kim function of depth $6$ and weight $251$.
@@ 5.2. A naive approach: computing the kernel of $\theta^{\#}_{d,v}$
[…]
@@
[…]
When $|S|=2$, by Table˜1, the minimal matrix size at which (5.2) holds is for $(d,v)=(14,65)$. When $d=14$, $296$ distinct Lyndon words show up in the image of $\theta^{\#}_{d}$. Thus, this brute force approach requires computing the kernel of a matrix of size $7354311\times 7370562$ over the ring $\mathbb{Q}[X_{1},\ldots,X_{296}]$, which is infeasible.
[…]
@@ 5.3. Non-existence of Chabauty–Kim functions in low degree
[…]
@@
[…]
Motivated by Proposition˜5.5, we have implemented the following algorithm in SageMath [sagemath]. It takes as input a tuple of integers $(s,d,v)$, where $s=|S|$, and outputs an integer $r$, which is the dimension of $\ker(M(\theta^{\#}_{d,v})(x))$ for a random $x\in\mathbb{Z}^{n}$. By Proposition˜5.5, if the output is 0, then there are no Chabauty–Kim functions in depth $d$ and degree $v$ for $|S|=s$. Moreover, if the output is $r$ for several iterations of the same tuple $(s,d,v)$, then it is likely that the dimension of $M(\theta^{\#}_{d,v})(x)$ is exactly $r$.
@@ 1. $(i)$
    Compute all monomials of degree $v$ in the variables $\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}$. These monomials form a basis of $(\mathcal{O}(U_{S})_{\leq d}[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}])_{v}$ as an $\mathcal{O}(U_{S})_{\leq d}$-module.
@@ 2. $(ii)$
    Compute all monomials of degree $v$ in the variables $\Phi_{\lambda}^{\rho}$ for $\rho\in\{\tau_{\ell},\sigma_{2k+1}\}$ and $\lambda=e_{0},e_{1},e_{1}e_{0}^{2k}$. These monomials form a basis of $(R_{\Phi,d})_{v}$ as an $\mathcal{O}(U_{S})_{\leq d}$-module.
@@ 3. $(iii)$
    Compute the set of Lyndon words $\mathcal{L}_{\geq-d}$ of $\Sigma$ to express $\mathcal{O}(U_{S})_{\leq d}$ as a polynomial algebra $\mathbb{Q}[\mathcal{L}_{\geq-d}]=\mathbb{Q}[X_{1},\ldots,X_{n}]$.
@@ 4. $(iv)$
    Compute $\theta_{d,v}^{\#}(\log^{\mathfrak{u}})$, $\theta_{d,v}^{\#}(\operatorname{Li}_{1}^{\mathfrak{u}}),\ldots,\theta_{d,v}^{\#}(\operatorname{Li}_{d}^{\mathfrak{u}})$ as elements of $\mathbb{Q}[X_{1},\ldots,X_{n}]$.  [^8] This step is the bottleneck of the algorithm, since it requires converting all words of length $d$ in $2$ letters into the dual PBWL basis.
@@ 5. $(v)$
    Choose a random $n$-tuple $x\in\mathbb{Z}^{n}$ and evaluate each of the polynomials $\theta_{d,v}^{\#}(\log^{\mathfrak{u}}))$, $\theta_{d,v}^{\#}(\operatorname{Li}_{i}^{\mathfrak{u}})$ at $x$.
@@ 6. $(vi)$
    Build the matrix $M(\theta_{d,v}^{\#})(x)$ with coefficients in $\mathbb{Q}$.
@@ 7. $(vii)$
The computation was carried out on The Ohio State University’s high performance cluster, and took around $66$ hours across $96$ cores, using 432432GB of memory. The overwhelming majority of this time was spent converting words in the shuffle algebra $\mathcal{O}(U_{S})$ into the dual PBWL basis (Poincaré–Birkhoff–Witt basis constructed on the Lyndon basis) using an improved version of SageMath’s built-in ShuffleAlgebra.to_dual_pbw_element method, rewritten to enable parallelisation, and to fix inefficiencies.
@@
Remark 5.10.
Originally, it was not clear to us that a motivic Chabauty–Kim function of depth $6$ and weight $18$ existed. Running the algorithm with $(2,6,18)$ made us suspect the existence of such a function. The elimination theory performed in Section˜6 gives a function of depth $6$ and degree $26$, which is a multiple of $(\log^{\mathrm{u}})^{6}$, and thus has a factor of degree $20$. The fact that the algorithm with $(2,6,18)$ returned $1$ made us look for an additional factor of degree $2$, which turned out to be $F^{|1|}_{2,2}$ defined by (4.2). In this sense, Algorithm 5.7 guided our search. Remark˜5.9 explains and quantifies why a brute force approach to finding a function of depth $6$ and degree $18$ is infeasible (contrary to the case $|S|=1$). Finally, we note that the $(2,6,18)$ calculation runs fairly quickly on a home computer ($10$ seconds), whereas the $(2,16,17)$ calculation took $66$ hours on a cluster.
@@ 6. The geometric step via resultants
Inspired by ideas in [DCJ], we introduce a method using resultants, whose goal is to produce a polylogarithmic motivic Chabauty–Kim function of depth $\leq d$, i.e., a non-zero element in the kernel of the homomorphism (Definition˜4.1)
@@
[…]
@@ 6.1. The equations defining the map $\theta^{\#}$
Making Definition˜4.1 explicit in the case $|S|=2$ gives the formulas:
@@ 6.2. The elimination process
We now describe the process of “eliminating” variables $\Phi_{\lambda}^{\rho}$, i.e., expressing them in terms of other variables $\Phi_{\lambda^{\prime}}^{\rho^{\prime}}$ and images under $\theta^{\#}$ in order to obtain a Chabauty–Kim function. The system of equations described in Section˜6.1 is linear with respect to the variables $\Phi^{\sigma_{2n-1}}_{e_{1}e_{0}^{2n-2}}$ and the variables $\Phi_{e_{1}}^{\tau_{p}}$ and $\Phi_{e_{1}}^{\tau_{q}}$. Consequently, we first seek to eliminate these variables. Eventually, the elimination of the variables $\Phi^{\tau_{p}}_{e_{0}}$ and $\Phi^{\tau_{q}}_{e_{0}}$ requires the use of a resultant. Throughout this section, we will be working with polynomials with coefficients in the fraction field $K=\mathrm{Frac}(\mathcal{O}(U_{S}))$.
@@ 6.2.2. Elimination of $\Phi_{e_{1}}^{\tau_{p}}$ and $\Phi_{e_{1}}^{\tau_{q}}$
[…]
@@
Since $F^{|1|}_{2,2}$ is of degree $2<18$, we deduce by Theorem˜5.8 that $\theta^{\#}(F^{|1|}_{2,2})\neq 0$ even though $F^{|1|}_{2,2}\in\mathcal{I}^{\operatorname{PL}}_{S^{\prime},2}$ when $|S^{\prime}|=1$. One can also compute this fact by hand.